This Question: 1 pt This Test: 50 pts possible Betore-tax cost of debt and after-tax cost of debt David Abbot i bank, and to repay the koan he will make 360 monthly payments (principal and interest) of $1,161.11 per month over the next 30 years. David can s buying a new house, and he is taking out a 30-year mortgage. David will borrow $209,000 from a payments on his mortgage from his taxable income, and a. What is the before-tax interest rate (per year) on Davids loan? b. What is the after-tax interest rate that David is paying? based on his income, David is in the 30% tax bracket a. The belore-tax interest rate (per year) on Davids loan is []% (Round to one decimal place ) b.The after tax interest rate that Dand is pay ng s % (Round to two de mal places ) Seudy
Answer #1

Answers:

a Before tax interest rate 5.3
b after tax interest rate 3.71

Calculations:

Here n = 30*12 = 360 (nper)

PV = 209,000

PMT = $1161.11

We have to find the rate. Let annual rate be x%. Thus monthly rate = x%/12

Thus using the PMT function in excel we get: PMT(X%/12, 360, 209,000) = 1161.11

Solving the above equation we get x = 5.3%

Thus before tax interest rate = 5.3%

After tax interest rate = 5.3%*(1-30%) = 3.71%